The complement of an angle measures 10 less than one-fifth of its supplement. Find the measure of the angle.
Answer by nerdybill(7384)

The complement of an angle measures 10 less than one-fifth of its supplement. Find the measure of the angle.
.
Let x = the angle
then
90-x = the complement
180-x = the supplement
.
90-x = (1/5)(180-x)
Multiplying both sides by 5:
450-5x = 180-x
450 = 180+4x
270 = 4x
67.5 degrees = x